Which of the following are equivalent to the statement AA and BB are independent?
A. P(A)P(B)=P(A and B)P(A)P(B)=P(A and B)
B. AA and BB cannot occur at the same time.
C. Knowing that AA occurs means BB cannot occur.
D. P(B|A)=P(B)P(B|A)=P(B)
E. P(A and B)=0P(A and B)=0
F. AA can happen without BB happening.
G. Knowing that AA occurs does not change the likelihood of BB
H. P(A|B)=P(A)P(A|B)=P(A)
Answer #1

Solution :

If A and B are independent then,

P(A and B) = P(A) P(B) ,

P(B | A) = P(B)

P(A | B) = P(A) Options A, D, and H are equivalent.